# From “Worse is Better” to Better: Lessons from a Mixed Methods Study of Ansible’s Challenges

### v1

@@ 3.3.3. Analysis
We developed five emergent codebooks using thematic analysis (Braun and Clarke, 2006; Lazar et al., 2017), following a bottom-up approach that allowed themes, codes to emerge inductively from our data. We started by reading the transcripts to highlight recurring ideas, creating initial codes, and iteratively refining them into coherent themes ( 3 and 4 in Figure 3). Our codebooks are related to the five key thematic areas as follows: (1) Background: Exploring why participants use Ansible and the specific contexts in which they apply it; (2) Advantages: Identifying the key motivations behind adoption; (3) Challenges: Documenting difficulties and limitations; (4) Recommendations: Capturing participants’ wishes; (5) Future: Addressing potential future developments.
@@ 4. RQ1: Common Issues in Ansible Usage
This section presents the challenges encountered by Ansible practitioners, obtained from the topic modeling analysis and interview data. We identified seven primary challenge categories presented in descending order of prevalence.
@@ 4.1.1. Topic Analysis
[…]
@@ 4.1.2. Interviews
Participants described various control flow issues. For instance, B1 mentioned frustration with sequential execution slowing runs and a lack of synchronization primitives (n = 2). In more advanced setups, complexities grew around managing dependencies (n = 3).Beyond these granular concerns, a handful of participants reflected on code structure/organization difficulties (n = 5). In contrast, others cited Ansible’s inherently stateless (n = 5) nature as a double-edged sword — simpler to run but lacking built-in features for preserving or comparing states. The push model/agentless (n = 5) approach was both praised and critiqued: praised for minimal overhead yet lamented for complicating continuous enforcement.
@@ 4.2.1. Topic Analysis
[…]
@@ 4.2.2. Interviews
Reliability issues emerged in discussions around debugging. Multiple participants reported that Ansible is hard to debug (n = 10) when playbooks fail mid-run. B2 described moments when “it would get stuck for 15 minutes in a supposedly 1 min script before I forced it to close without any visible changes on the terminal.”. Some lamented the lack of infrastructure debugging tools (n = 5) built into Ansible, while others pointed to bad error messages when modules returned cryptic outputs. As B4 put it, “I actually believe that there’s not a proper way to debugging this [Infrastructure as Code] type of scripts and tools.”.
@@ 4.3.1. Topic Analysis
[…]
@@ 4.3.2. Interviews
The theme lack of support touched on problems such as missing features that require modules (n = 3), limited OS support (n = 4), and Python version constraints (n = 1). Other participants criticized the lack of support for certain modules (n = 6). As E4 put it “There’s a lack of support for some modules (…). You often have to write your own or hope someone did it.”.
@@ 4.4.1. Topic Analysis
[…]
@@ 4.4.2. Interviews
Ansible’s usability emerged as a significant issue for many participants. It includes user onboarding issues and complaints about complex syntax. Participants expressed that it can be challenging for beginners (n = 10) to grasp the tool’s declarative nature, in particular for those that lacked a coding background (n = 2). Others felt they had to make a mental shift (n = 2), especially if they were used to more imperative scripting, and sometimes it can be hard to pick the right approach (n = 4). Participants also complained about complex syntax (n = 10). Ansible’s reliance on YAML and Jinja introduced the problems of YAML and Jinja complexity (n = 5 and 6). As E5 said, “I once spent 8 hours of work figuring out the escaping on a line of SQL because YAML is so bad at figuring out quotes (…) The hell that is escaping in YAML will never die.”.
@@ 4.5.1. Topic Analysis
[…]
@@ 4.5.2. Interviews
Scalability limitations (n = 8) were frequently mentioned by participants managing large infrastructures. B1 stated: “When you run a playbook for like 100 hosts, it gets really confusing trying to make sense of the messages.”. Participants specifically described problems related to managing inventories (n = 1) and code awareness (n = 4) in large projects. A related source of concern was slow execution (n = 7). While some liked the simplicity of SSH-based workflows, a few individuals cited agentless (n = 4) as detrimental in high-volume scenarios. E4 was particularly direct: “It does not scale when you start to manage 100th of servers, it becomes slow because there is no agent.”.
@@ 4.6.2. Interviews
Participants noted that certain aspects of Ansible maintenance remain nontrivial. In particular, several mentioned that removing configurations is difficult (n = 3) — as modifying the playbook alone does not guarantee that older changes get undone without explicit tasks. E2 explained that “Doing things like removing config from hosts tends not to work so well in Ansible,”. Configuration drift (n = 5) was also mentioned. As I3 described “Deletes are hard: one cannot merely delete parts of the Ansible playbook and expect those changes to be undone in hosts. Instead, we must change or even add tasks undoing the configuration we want gone.”. Other participants mentioned that Ansible can be hard to maintain (n = 4) and were concerned about deprecations (n = 4).
@@ 4.7.1. Interviews
Another concern revolved around information gaps, particularly documentation that is either overwhelming or incomplete. E2 talked about misleading information: “Another issue is incorrect or misleading answers on forums and newsgroups. Beginners may struggle to distinguish good advice from bad.”. The theme of documentation gaps (n = 8) appeared in several responses, with participants describing the difficulty of locating relevant or clear information among Ansible’s extensive (but sometimes inconsistent) references. Although multiple participants praised Ansible’s core documentation (described in Section 5), they also identified a steep learning curve for interpreting specific advanced.
@@ 5. RQ2: Factors Influencing Ansible Adoption
This section analyzes how interview participants first encountered Ansible and their reasons for selecting it over alternative IaC solutions. We go over the thematic areas Background and Advantages described in Section 3.3.3.
@@
Initial Exposure to Ansible. Participants described how they discovered Ansible. The most common way was in professional settings (n = 5), where it was already an established tool. Social media platforms also played a notable role in exposure (n = 4), particularly X. As E5 recounted: “In 2016, [someone] posted a tweet about a new configuration management tool (…) I had a look.” Some participants also independently sought out Ansible (n = 3).
@@
Accessibility & Usability. A significant theme in our codebook is the role of Ansible’s accessibility and usability in driving adoption. Several participants described it as Easy to use (n = 12), Easy to onboard to new people (n = 5), and with a Low Barrier to Entry for Setup (n = 6) highlighting how they only needed to master a limited set of concepts—such as basic playbook structures and inventory configuration—before managing a wide range of environments. B1 remarked, “To get things done, I only really needed to understand a small set of concepts regarding how Ansible works (…) ”. Participants also cited Ansible’s Readable syntax (n = 5) and Good Logs (n = 1). E6 highlighted that “[we] were looking into Ansible because it has a nice YAML syntax”.
@@
Tool Qualities and Ecossystem. Beyond accessibility, Ansible was praised for its tool qualities. Participants highlighted its Efficient use of resources, with one participant citing Lightweight resource consumption and three mentioning Scalability. Efficiency was also linked to Cost-Effectiveness, with B4 saying, “Ansible is like a [inexpensive] version [of Kubernetes] for scalability”. Another advantage of Ansible is its Flexibility, since it can be a One-Tool Solution (n = 2) capable of managing heterogeneous environments (e.g., Windows, Linux, or network devices). As E3 puts it, the “One tool that does it all”. Several participants emphasized that Ansible’s Compatibility (n = 4) with various operational requirements and existing workflows set it apart from other tools, and its Agentless (n = 9) model that depends solely on SSH. Idempotence (n = 2) was also considered important for participants, ensuring the same playbook can be applied repeatedly to converge infrastructure into a consistent state. As I3 described: “Some commands are dangerous to invoke twice.” From an Ecosystem standpoint, respondents highlighted features such as Easy to extend (n = 3), Good Tooling (n = 1), and the fact that It’s implemented in Python (n = 2).
@@
Participants frequently framed their rationale in comparative terms, specifically highlighting Ansible’s advantages over alternatives like Puppet, Chef, and Salt. A common comparative advantage cited was the elimination of agent management overhead. E5 provided an illustrative anecdote highlighting Ansible’s comparative operational model advantages: “(…) a place I was working was using Puppet, and the admin was woken up at 3 am from a page, fixed the issue, went to bed, and then Puppet restored the broken configuration, so he woke up again, fixed the Puppet, and then went to bed. Then, was woken up again because Puppet updated itself from git.”.
@@
Community. A third central theme centers on Ansible’s Community. The tool’s Popularity (n = 4) and Community (n = 3) are significant advantages. E1 noted: “Ansible’s advantages (…) are it being a well-supported IaC tool. Since there is a large community around it, it is easier to find solutions to specific problems online.” I1 summarized it as: “Ansible was just the hot topic at the time.” Participants also mentioned that Ansible’s strong market presence often led their organizations or academic institutions to mandate its use (n = 7) (“it was standard by the time I joined the team”, I3). Moreover, if it was not mandatory, it was often Recommended (n = 2) by colleagues or professors.
@@ 6. RQ3: Improving Ansible
This section analyzes what our interviewees would like to see in Ansible and how they see its future. We go over the thematic areas Future and Recommendations described in Section 3.3.3.
6.1. Future of Ansible
When asked about Ansible’s future trajectory, the participants provided various assessments reflecting optimism and caution. A small group (n = 2) expressed expressing concerns that Ansible might become less relevant over time “As more compute ends up in serverless and ephemeral hosts on public and private cloud, Ansible is becoming less relevant.” (E2). The majority of participants took a more Neutral stance, will likely remain a practical choice for specific use cases (n = 5) — as I3 put it “My expectation is that future usage of Ansible will tend towards maintaining legacy systems, and use as a configuration-specific scripting language ”. Others see Ansible retaining its current role (n = 4) in the ecosystem. Other respondents held a Positive outlook, Hoping that the tool will expand in scope and capabilities (n = 2) like described by E5, “I honestly hope that everyone uses. I am teaching my 5-year-old to program, and we’re talking about using Ansible to update his streaming service (mirroring RSS feeds from YouTube) and updating his FDroid repository! I hope that people get empowered to build their systems .”. Others noted that Demand is growing (n = 3) among organizations and peers alike. E2 summarizes their answer by saying “I think the general gist of my opinion is: Ansible is another ‘Worse Is Better’ system. That’s both great and annoying in equal measure, and is why Ansible is as popular as it is.” [^1] This quote inspired the title of our paper..
6.2. Suggestions and Improvements
@@
While participants generally acknowledged Ansible’s strengths, they also offered concrete recommendations to improve.
@@
Documentation and Learning Resources. A recurring concern expressed by participants (n = 8) was the need to improve documentation. E4 described the current official documentation as scattered, noting that it “is poor considering the other configuration tools (take a look at Puppet’s). The websites are not appealing and not properly organized. Strong concepts are scattered over different pages.”. I3 identified specific documentation challenges regarding module organization: “Documentation needs a big revamp: it’s not clear what modules are built-in or part of Ansible Galaxy.”. The participants also called for better onboarding resources (n = 5) and more education (n = 7). As E3 put it “ When you educate beginners, they can identify incorrect information and avoid it.”.
@@
Debugging and Troubleshooting. Several participants highlighted limitations in Ansible’s current debugging capabilities. Better debugging tools (n = 5) and improved error messages (n = 6) were frequently mentioned. E2 stressed that “error messages when plugins fail could be improved,” as they can sometimes be “misleading”. I2 reinforced this point, saying that “more specific error messages would be very helpful. Understanding exactly where an issue presents is the most important part of the debugging process.”. Another concept frequently raised was the lack of rollback-on-failure support (n = 4).
@@
Performance and Scalability. Participant also critiqued performance and mentioned the need to improve speed (n = 6) and improve parallelization (n = 4). As E2 put “Slow runs is an eternal pain point.”. E1 advocated for integrating solutions akin to Mitogen (Mitogen Contributors, 2025), to help optimize remote execution over SSH to reduce overhead and B1 proposed a tag-based approach: “Having the ability to set group tags for tasks, and have groups of tasks be able to run in parallel would be a decent enough solution (tasks with tag A can run in parallel with tasks with tag B).”.
@@
Advanced Programming Capabilities. A group of participants called for the expansion of Ansible’s programming model. While some participants advocated for more high-level language concepts (n = 3) — including real loops, complex conditionals, and synchronization primitives — others (n = 4) stressed state management. E7, for instance, argued for “functionality in Ansible to handle state management effectively” and provided an extensive analysis of Ansible’s state management limitations, saying, “even though Ansible claims to be completely stateless, which is supposed to be its strength, I’m not so convinced. Without state management, it’s challenging to handle certain problems.”.
@@
Participant Experience Levels. While some recommendations, such as providing more education, were popular across experience levels, other priorities varied according to participants’ familiarity with Ansible. Beginners often highlighted the need for better onboarding (n = 2) and rollback on failure (n = 2). In contrast, Experienced users (n = 4) tended to focus on improving speed and other performance enhancements, reflecting the operational demands of managing more extensive infrastructures.

### v2

@@ 3.3.3. Analysis
We developed five emergent codebooks using thematic analysis (Braun and Clarke, 2006; Lazar et al., 2017), following a bottom-up approach that allowed themes and codes to emerge inductively from our data. We started by reading the transcripts to highlight recurring ideas, creating initial codes, and iteratively refining them into coherent themes ( 3 and 4 in Figure 2). Our codebooks are related to the five key thematic areas as follows: (1) Background: Exploring why participants use Ansible and the specific contexts in which they apply it; (2) Advantages: Identifying the key motivations behind adoption; (3) Challenges: Documenting difficulties and limitations; (4) Recommendations: Capturing participants’ wishes; (5) Future: Addressing potential future developments.
@@ 4. RQ1: Common Issues in Ansible Usage
This section presents the seven primary challenge categories encountered by Ansible practitioners, obtained from the topic modeling analysis and interview data, in descending order of prevalence.
@@ 4.1.1. Topic Analysis
The number of references to Loops in posts increased from 2.9% of all posts in 2013 to 7.4% in 2022, before decreasing to 4.3% in 2024 ( ). Contrarily, the number of posts related to Variable Management decreased from 8.9% in 2013 to 5.4% in 2024 ( ). This evolution suggests that, while loop issues persist and even increase, Ansible may be making progress in addressing the challenges related to management of variables.
@@ 4.1.2. Interviews
Participants described various control flow issues. For instance, B1 mentioned frustration with sequential execution slowing runs and a lack of synchronization primitives (n = 2). In more advanced setups, complexities grew around managing dependencies (n = 3). Beyond these granular concerns, a handful of participants reflected on code structure/organization difficulties (n = 5). In contrast, others cited Ansible’s inherently stateless (n = 5) nature as a double-edged sword — simpler to run but lacking built-in features for preserving or comparing states. The push model/agentless (n = 5) approach was both praised and criticized: praised for minimal overhead yet lamented for complicating continuous enforcement.
@@ 4.2.1. Topic Analysis
In 2013, Permission Issues represented just 1.4% of the total posts in that year, but by 2024, this figure had increased to 3.2% ( ). This increase may suggest that Ansible practitioners are encountering more permission-related issues, likely due to interactions with hosts that have more strict security configurations.
@@ 4.2.2. Interviews
Reliability issues emerged in discussions around debugging. Multiple participants reported that Ansible is hard to debug (n = 10). I7 tries to avoid Ansible to debug complex cases: “if there’s some more complex stages in my playbooks what I try to do is to transfer those to a shell script that I can pass outside of Ansible and I can test those separately.”. B2 described moments when “it would get stuck for 15 minutes in a supposedly 1 min script before I forced it to close without any visible changes on the terminal.”. Some lamented the lack of infrastructure debugging tools (n = 5) built into Ansible, while others pointed to bad error messages when modules returned cryptic outputs. As B4 put it, “I actually believe that there’s not a proper way to debugging this [Infrastructure as Code] type of scripts and tools.”.
@@ 4.3.1. Topic Analysis
The number of posts related to AWX, launched in 2017, has increased significantly from 1.6% of the overall posts in its release year to 9.8% in 2024 ( ). Another topic that saw significant growth was Integration Issues, increasing from 0.8% in 2013 to 5.1% in 2024 ( ). This suggests that as Ansible matured, practitioners increasingly sought solutions for integrating with external services. In the opposite direction, posts related to Module Development decreased from 1.6% in 2013 to 0.6% in 2024 ( ).
@@ 4.3.2. Interviews
The theme lack of support touched on problems such as missing features that require modules (n = 3), limited OS support (n = 4), and Python version constraints (n = 1). Other participants criticized the lack of support for certain modules (n = 7). As E4 put it “There’s a lack of support for some modules (…). You often have to write your own or hope someone did it.”.
@@ 4.4.1. Topic Analysis
In 2013, issues related to Jinja2 were not as prominent, only 2.4% of the posts that year. However, Jinja2-related issues peaked at 6.9% in 2021, before decreasing to 3.8% by 2024 ( ).
@@ 4.4.2. Interviews
Ansible’s usability emerged as a significant issue for many participants. It includes user onboarding issues and complaints about complex syntax. Participants expressed that it can be challenging for beginners (n = 12) to grasp the tool’s declarative nature, in particular for those that lacked a coding background (n = 2). Others felt they had to make a mental shift (n = 4), especially if they were used to more imperative scripting. As E8 said, “If you’ve never seen Terraform, CloudFormation (…) it’ll look weird and alienating.”. Moreover, sometimes it can be hard to pick the right approach (n = 4). Participants also complained about complex syntax in general (n = 4). Ansible’s reliance on YAML and Jinja introduced the problems of YAML and Jinja complexity (n = 5 and 6). As E5 said, “I once spent 8 hours of work figuring out the escaping on a line of SQL because YAML is so bad at figuring out quotes (…) The hell that is escaping in YAML will never die.”.
@@ 4.5.1. Topic Analysis
Inventory management issues have remained relatively stable between 2013 and 2024, consistently accounting for between 1.8% and 2.5% of the total yearly posts ( ).
@@ 4.5.2. Interviews
Scalability limitations (7) were frequently mentioned by participants managing large infrastructures. B1 stated: “When you run a playbook for like 100 hosts, it gets really confusing trying to make sense of the messages.”. Participants specifically described problems related to managing inventories (n = 1) and code awareness (n = 4) in large projects. A related source of concern was slow execution (n = 8). While some liked the simplicity of SSH-based workflows, a few individuals cited agentless (n = 4) as detrimental in high-volume scenarios. E4 was particularly direct: “It does not scale when you start to manage 100th of servers, it becomes slow because there is no agent.”.
@@ 4.6.2. Interviews
Participants noted that certain aspects of Ansible maintenance remain nontrivial. In particular, several mentioned that removing configurations is difficult (n = 3) — as modifying the playbook alone does not guarantee that older changes get undone without explicit tasks. E2 explained that “Doing things like removing config from hosts tends not to work so well in Ansible,”. Configuration drift (n = 5) was also mentioned. As I3 described “Deletes are hard: one cannot merely delete parts of the Ansible playbook and expect those changes to be undone in hosts. Instead, we must change or even add tasks undoing the configuration we want gone.”. Other participants mentioned that Ansible can be hard to maintain (n = 5) and were concerned about deprecations (n = 5).
@@ 4.7.1. Interviews
Another concern revolved around information gaps, particularly documentation that is either overwhelming or incomplete. E2 talked about misleading information: “Another issue is incorrect or misleading answers on forums and newsgroups. Beginners may struggle to distinguish good advice from bad.”. The theme of documentation gaps (n = 9) appeared in several responses, with participants describing the difficulty of locating relevant or clear information among Ansible’s extensive (but sometimes inconsistent) references. Although multiple participants praised Ansible’s core documentation (described in Section 5), they also identified a steep learning curve for interpreting specific advanced.
@@ 5. RQ2: Factors Influencing Ansible Adoption
This section analyzes how interview participants first encountered Ansible and their reasons for selecting it over other IaC solutions. We go over Background and Advantages described in Section 3.3.3.
@@
Initial Exposure to Ansible. Participants described how they discovered Ansible. The most common way was in professional settings (n = 6), where it was already an established tool. Social media platforms also played a notable role in exposure (n = 4), particularly X. As E5 recounted: “In 2016, [someone] posted a tweet about a new configuration management tool (…) I had a look.” Some participants also independently sought out Ansible (n = 3).
@@
Accessibility & Usability. A significant theme in our codebook is the role of Ansible’s accessibility and usability in driving adoption. Several participants described it as Easy to use (n = 13), Easy to onboard to new people (n = 7), and with a Low Barrier to Entry for Setup (n = 6) highlighting how they only needed to master a limited set of concepts—such as basic playbook structures and inventory configuration—before managing a wide range of environments. B1 remarked, “To get things done, I only really needed to understand a small set of concepts regarding how Ansible works (…) ”. Participants also cited Ansible’s Readable syntax (n = 6) and Good Logs (n = 1). E6 highlighted that “[we] were looking into Ansible because it has a nice YAML syntax”.
@@
Tool Qualities and Ecossystem. Beyond accessibility, Ansible was praised for its tool qualities. Participants highlighted its Efficient use of resources, with one participant citing Lightweight resource consumption and three mentioning Scalability. Efficiency was also linked to Cost-Effectiveness, with B4 saying, “Ansible is like a [inexpensive] version [of Kubernetes] for scalability”. Another advantage of Ansible is its Flexibility, since it can be a One-Tool Solution (n = 2) capable of managing heterogeneous environments (e.g., Windows, Linux, or network devices). As E3 puts it, the “One tool that does it all”. Several participants emphasized that Ansible’s Compatibility (n = 6) with various operational requirements and existing workflows set it apart from other tools, and its Agentless (n = 9) model that depends solely on SSH. Idempotence (n = 2) was also considered important for participants, ensuring the same playbook can be applied repeatedly to converge infrastructure into a consistent state. As I3 described: “Some commands are dangerous to invoke twice.” From an Ecosystem standpoint, respondents highlighted features such as Easy to extend (n = 4), Good Tooling (n = 1), and the fact that It’s implemented in Python (n = 2).
@@
Participants frequently framed their rationale in comparative terms, highlighting Ansible’s advantages over alternatives like Puppet, Chef, and Salt. A common comparative advantage cited was the elimination of agent management overhead. E5 provided an illustrative anecdote highlighting Ansible’s comparative operational model advantages: “(…) a place I was working was using Puppet, and the admin was woken up at 3 am from a page, fixed the issue, went to bed, and then Puppet restored the broken configuration, so he woke up again, fixed the Puppet, and then went to bed. Then, was woken up again because Puppet updated itself from git.”.
@@
Community. A third central theme centers on Ansible’s Community. The tool’s Popularity (n = 4) and Community (n = 4) are significant advantages. E1 noted: “Ansible’s advantages (…) are it being a well-supported IaC tool. Since there is a large community around it, it is easier to find solutions to specific problems online.” I1 summarized it as: “Ansible was just the hot topic at the time.” Participants also mentioned that Ansible’s strong market presence often led their organizations or academic institutions to mandate its use (n = 8) (“it was standard by the time I joined the team”, I3). Moreover, if it was not mandatory, it was often Recommended (n = 3) by colleagues or professors.
@@ 6. RQ3: Improving Ansible
This section describes the Recommendations described in Section 3.3.3. While participants generally acknowledged Ansible’s strengths, they also offered concrete recommendations to improve.
@@
Documentation and Learning Resources. A recurring concern expressed by participants (n = 10) was the need to improve documentation. E4 described the current official documentation as scattered, noting that it “is poor considering the other configuration tools (take a look at Puppet’s). The websites are not appealing and not properly organized. Strong concepts are scattered over different pages.”. I3 identified specific documentation challenges regarding module organization: “Documentation needs a big revamp: it’s not clear what modules are built-in or part of Ansible Galaxy.”. E8 found the documentation examples helpful but noted that “sometimes they don’t [provide] all the examples” The participants also called for better onboarding resources (n = 6) and more education (n = 9). As E3 put it “ When you educate beginners, they can identify incorrect information and avoid it.”.
@@
Debugging and Troubleshooting. Several participants highlighted limitations in Ansible’s current debugging capabilities. Better debugging tools (n = 5) and improved error messages (n = 6) were frequently mentioned. E2 stressed that “error messages when plugins fail could be improved,” as they can sometimes be “misleading”. I2 reinforced this point, saying that “more specific error messages would be very helpful. Understanding exactly where an issue presents is the most important part of the debugging process.”. Another concept frequently raised was the lack of rollback-on-failure support (n = 5).
@@
Performance and Scalability. Participant also critiqued performance and mentioned the need to improve speed (n = 7) and improve parallelization (n = 4). As E2 put “Slow runs is an eternal pain point.”. E1 advocated for integrating solutions akin to Mitogen (Mitogen Contributors, 2025), to help optimize remote execution over SSH to reduce overhead and B1 proposed a tag-based approach: “Having the ability to set group tags for tasks, and have groups of tasks be able to run in parallel would be a decent enough solution (tasks with tag A can run in parallel with tasks with tag B).”.
@@
Advanced Programming Capabilities. A group of participants called for the expansion of Ansible’s programming model. While some participants advocated for more high-level language concepts (n = 3) — including real loops, complex conditionals, and synchronization primitives — others (n = 5) stressed state management. E7, for instance, argued for “functionality in Ansible to handle state management effectively” and provided an extensive analysis of Ansible’s state management limitations, saying, “even though Ansible claims to be completely stateless, which is supposed to be its strength, I’m not so convinced. Without state management, it’s challenging to handle certain problems.”.
@@
Participant Experience Levels. While some recommendations, such as providing more education, were popular across experience levels, other priorities varied according to participants’ familiarity with Ansible. Beginners often highlighted the need for better onboarding (n = 2) and rollback on failure (n = 3). In contrast, Experienced users (n = 4) tended to focus on improving speed and other performance enhancements, reflecting the operational demands of managing more extensive infrastructures.